The school that Micaela goes to is selling tickets to a spring musical. On the first day of ticket sales the school sold 11 adult tickets and 9 student tickets for a total of $138. The school took in $118 on the second day by selling 1 adult ticket and 14 student tickets. Find the price of an adult ticket and the price of a student ticket.
Answer by ankor@dixie-net.com(22740)

The school that Micaela goes to is selling tickets to a spring musical.
On the first day of ticket sales the school sold 11 adult tickets and 9 student tickets for a total of $138.
The school took in $118 on the second day by selling 1 adult ticket and 14 student tickets.
Find the price of an adult ticket and the price of a student ticket.
:
let a = the price of adult tickets
let s = the price of student's
:
Write an equation for each day
11a + 9s = 138.00
and
1a + 14s = 118.00
a = (118-14s); use this form for substitution
:
11(118-14s) + 9s = 138
1298 - 154s + 9s = 138
-154s + 9s = 138 - 1298
-145s = -1160
s = -1160/-145
s = $8.00 is the student's price
Find a
a = 118 - 14(8)
a = 118 - 112
a = $6.00 is the adult price
:
Check this using the 1st equation, replace a and s
11a + 9s = 138
11(6) + 9(8) =
66 + 72 = 138